# Constrained Dynamics for Searching Saddle Points on General Riemannian Manifolds

### v1

@@ 1 Introduction
In this work, we develop a constrained saddle dynamics on general Riemannian manifolds and analyze the theoretical properties of the continuous dynamics and the resulting discretized algorithms. We further demonstrate the effectiveness of algorithms through electronic excited-state calculations on a standard benchmark molecular system.
@@ 1.2 Contributions
Finally, we demonstrate the effectiveness of the developed algorithm on linear eigenvalue problems (Section 5.1) and electronic excited-state calculations for a standard benchmark molecule (Section 5.2). We also corroborate numerically the influence of problem data and the importance of removing parametrization redundancies when searching for SPs.
@@ 5.1 Linear eigenvalue problem
¯¯¯from jax import random, numpy
¯¯¯key = random.PRNGKey(randseed) # random seed
¯¯¯key1, key2 = random.split(key)
¯¯¯X, _ = numpy.linalg.qr(random.normal(key1, (n, p)))
¯¯¯P = X @ X.T
¯¯¯Gamma = random.normal(key2, (n, n))
¯¯¯Gamma = (Gamma + Gamma.T) / 2.0
¯¯¯Gamma = proj_tangent(P, Gamma) # project onto the tangent space
¯¯¯V = Gamma @ X # horizontal lift
@@
¯¯¯key = random.PRNGKey(randseed) # random seed
¯¯¯key1, key2 = random.split(key)
¯¯¯X, _ = numpy.linalg.qr(X_ref + beta * random.normal(key1, (n, p)))
¯¯¯# repeat previous procedures to create P, Gamma, and V
Here $X_{\rm ref}$ is either the global minimizer or index-1 constrained SP. The empirical success rates of the two algorithms at different perturbation levels are recorded in Table 1.
@@
It turns out that the Grassmann version converges surprisingly well across the board to the index-1 constrained SP, whereas the Stiefel one often gets attracted by the global minimizer, with an empirical success rate of approximately 30% if randomly initialized. The rate can be tremendously improved if the initial point is selected around the desired index-1 constrained SP. We only use the Grassmann version for the subsequent experiments.
@@ 5.2 Electronic excited-state calculations
The core of electronic calculations for molecular systems is the electronic Schrödinger equation (ESE) [schrodinger1926undulatory], which is in fact a linear eigenvalue problem. Nevertheless, the ESE is intractable in general due to the curse of dimensionality. For the numerical purpose, various approximations have been proposed in some atomic-orbital basis, such as the full configuration interaction (FCI), Hartree-Fock (HF) methods, and post-HF methods [helgaker2000molecular], among others. Excited states define the optical and reaction properties of atoms and molecules [balzani2014photochemistry, mai2020molecular, turro2009principles]. Characterizing the excited states is challenging due to electron correlation effects [gonzalez2012progress]. Finding the constrained SPs of the quantum chemical approximated methods in use arises as a natural methodology [burton2022energy, burton2020energy, cances2006computing, lewin2004solutions, marie2023excited, saade2024excited]. Moreover, these approximations usually come together with manifold structures. In the following, we briefly introduce the HF methods and report the numerical results of finding constrained SPs as candidates of excited states, with FCI calculations (performed by PySCF [sun2020recent]) as reference. More advanced levels of theory, such as the complete active space self-consistent field method [roos1980complete], involve complicated manifolds and will be investigated in a parallel work. Before proceeding, we remark that our methodology falls into the class of state-specific methods in quantum chemistry for electronic excited states; other popular ones include linear response theory [cances1999time, casida1995time, grazioli2026critical, olsen1985linear] and state-average methods for multiconfigurational approximations [werner1981quadratically].
@@
Implementation details. We consider the H2 molecule with its bond length varied from 0.1 a.u. to 4.0 a.u.222“a.u.” is an abbreviation of “atomic unit” for various measurements, which is “Bohr” for length and “Hartree” for energy. 1 Bohr $\approx$ $5.29\times\texttimes 10^{-11}$ m., 1 Hartree $\approx 4.36\times 10^{-18}$ J., with a spacing of 0.1 a.u. The molecule is described by RHF with the 6-31G basis set ($N_{\rm elec}=2$, $N_{{\rm b}}=4$, $N_{{\rm o}}=1$) [ditchfield1971self, hehre1972self], or in the chemical notation, RHF/6-31G. The number of total degrees of freedom (DOFs) is thus $N_{{\rm o}}(N_{{\rm b}}-N_{{\rm o}})=3$. We investigate the energy landscape of RHF for the H2 molecule with the varying bond length by searching for the constrained SPs of indices $0\sim 3$. They are found by running Algorithm 2 from 1,000 random initial points (random seed $0\sim 999$); see the pseudocodes in Section 5.1 for initialization. For retraction, Algorithm 2 is equipped with the exponential mapping for the position part; the treatment for the direction part is similar to that in the previous subsection. The step size is specified as $\eta=10^{-1}$, which is not necessarily optimal. The maximum iteration number and convergence tolerance are respectively set as $maxit=\infty$ and $tol=10^{-6}$. The results are compared with those obtained by solving FCI/6-31G based on the RHF calculations, which is exact under the basis in use. In this setting, FCI gives the ground state and 15 excited states for the molecule. We remark that FCI calculations are unaffordable in general cases, since it involves $\mathcal{O}(\binom{N_{\rm b}}{N_{\rm elec}})$ DOFs.
@@
Figure 5: An overview of the constrained SPs of RHF found across the bond length interval on the H2 molecule. The blue solid and dashed lines represent the energies of FCI states belonging to the irreps $\rm A_{g}$} and $\rm A_{u}$}, respectively. The red dots, deepblue squares, yellow stars, and purple diamonds stand for the energies of the RHF SPs of indices 0, 1, 2, and 3, respectively.
@@
Figure 6: Zoom-in views of the constrained SPs of RHF. Left: an RHF index-1 SP disappears in (0.5, 0.6). Middle: an RHF index-2 SP emerges in (1.6, 1.7). Right: an RHF index-0 SP emerges in (2.6, 2.7).
@@
Results on the H2 molecule. An overview of the constrained SPs of RHF identified across the considered bond length interval is shown in Figure 5, together with the ground-/excited-state energies of FCI as a reference. It can be seen that the RHF energy landscape varies smoothly with the bond length. Note that FCI states are classified by their irreducible representations (irreps); for the H2 molecule in the computational point group $\rm D_{2h}$}, the relevant irreps are $\rm A_{g}$} and $\rm A_{u}$}, with eight FCI states belonging to each. It is observed that the RHF SPs are only able to describe a small subset (three or four) of FCI states in terms of their energies. This behavior is consistent with the fact that RHF neglects electronic correlation [barca2014communication]. The deficiency can be mitigated to some extent by resorting to post-HF methods.
@@
Our results also reveal numerically that the varying bond length, as an external parameter, can lead to disappearance or emergence of constrained SPs. Zoom-in views are given in Figure 6. Concretely, in the interval (0.5, 0.6), an index-1 SP gets close to the index-2 SP, in terms of energy, and disappears; in the interval (1.6, 1.7), an index-2 SP close to the index-3 SP emerges; and in the interval (2.6, 2.7), an index-0 SP (i.e., a local minimizer) close to the index-1 SP emerges.
@@
The above results provide proof-of-concept evidence for the effectiveness of our algorithms in excited-state calculations. Nonetheless, we shall point out that a comprehensive quantum chemical analysis of the obtained SPs, though beyond the scope of present work, is essential for practical applications. As a nonlinear approximation to the exact theory, RHF may yield critical points that lack physical meaning; e.g., a spurious non-global local minimizer emerges when the bond length exceeds 2.7 in the right panel of Figure 6. In addition, since RHF constitutes a low-dimensional approximation, a one-to-one correspondence between RHF SPs and excited states in the same energetic order no longer holds. It is also of great importance to develop schemes capable of navigating the nonconvex landscape efficiently [xu2025general], instead of random multi-start.
@@ Acknowledgements
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement EMC2 No 810367). The authors are grateful to Eric Cancès, Tony Lelièvre, and Panos Parpas for useful discussions.

### v2

@@ 1 Introduction
In this work, we develop a constrained saddle dynamics on general Riemannian manifolds and analyze the theoretical properties of the continuous dynamics and the resulting discretized algorithms. We further demonstrate the effectiveness of algorithms through electronic excited-state calculations on standard benchmark molecular systems.
@@ 1.2 Contributions
Finally, we demonstrate the effectiveness of the developed algorithm on linear eigenvalue problems (Section 5.1) and electronic excited-state calculations for standard benchmark molecules (Section 5.2). We also corroborate numerically the influence of problem data and the importance of removing parametrization redundancies when searching for SPs.
@@ 5.1 Linear eigenvalue problem
from jax import random, numpy
    key = random.PRNGKey(randseed) # random seed
    key1, key2 = random.split(key)
    X, _ = numpy.linalg.qr(random.normal(key1, (n, p)))
    P = X @ X.T
    Gamma = random.normal(key2, (n, n))
    Gamma = (Gamma + Gamma.T) / 2.0
    Gamma = proj_tangent(P, Gamma) # project onto the tangent space
    V = Gamma @ X # horizontal lift
@@
key = random.PRNGKey(randseed) # random seed
    key1, key2 = random.split(key)
    X, _ = numpy.linalg.qr(X_ref + beta * random.normal(key1, (n, p)))
    # repeat previous procedures to create P, Gamma, and V
Here $X_{\rm ref}$ is either the global minimizer or index-1 constrained SP. The empirical success rates of the two algorithms at different perturbation levels are recorded in Table 1. It turns out that the Grassmann version converges surprisingly well across the board to the index-1 constrained SP, whereas the Stiefel one often gets attracted by the global minimizer, with an empirical success rate of approximately 30% if randomly initialized. The rate can be tremendously improved if the initial point is selected around the desired index-1 constrained SP. We only use the Grassmann version for the subsequent experiments.
@@ 5.2 Electronic excited-state calculations
The core of electronic calculations for molecular systems is the electronic Schrödinger equation (ESE) [schrodinger1926undulatory], which is in fact a linear eigenvalue problem. Nevertheless, the ESE is intractable in general due to the curse of dimensionality. For the numerical purpose, various approximations have been proposed in some atomic-orbital basis, such as the full configuration interaction (FCI), Hartree-Fock (HF) methods, and post-HF methods [helgaker2000molecular], among others. Excited states define the optical and reaction properties of atoms and molecules [balzani2014photochemistry, mai2020molecular, turro2009principles]. Characterizing the excited states is challenging due to electron correlation effects [gonzalez2012progress]. Finding the constrained SPs of the quantum chemical approximated methods in use arises as a natural methodology [burton2022energy, burton2020energy, cances2006computing, lewin2004solutions, marie2023excited, saade2024excited]. Moreover, these approximations usually come together with manifold structures. In the following, we briefly introduce the HF methods and report the numerical results of finding constrained SPs as candidates of excited states, with FCI calculations (performed by PySCF [sun2020recent]) as reference. More advanced levels of theory, such as the complete active space self-consistent field method [roos1980complete], involve complicated manifolds and are investigated in a parallel work. Before proceeding, we remark that our methodology falls into the class of state-specific methods in quantum chemistry for electronic excited states; other popular ones include linear response theory [cances1999time, casida1995time, grazioli2026critical, olsen1985linear] and state-average methods for multiconfigurational approximations [werner1981quadratically].
@@
Implementation details. We consider the H2 and LiH molecules. The bond length of the H2 molecule is varied from 0.1 bohr2221 bohr $\approx$ $5.29\times\texttimes 10^{-11}$ m. to 4.0 bohr with a spacing of 0.1 bohr, while that of the LiH molecule is from 2.0 bohr to 6.0 bohr with a spacing of 0.2 bohr The H2 and LiH molecules are described by RHF with the 6-31G basis set ($N_{\rm elec}=2$, $N_{{\rm b}}=4$, $N_{{\rm o}}=1$) [ditchfield1971self, hehre1972self] and the STO-3G basis set ($N_{\rm elec}=4$, $N_{{\rm b}}=6$, $N_{\rm o}=2$) [hehre1969self], respectively. The numbers of total degrees of freedom (DOFs) are thus $N_{{\rm o}}(N_{{\rm b}}-N_{{\rm o}})=3$ for the H2 molecule and 8 for the LiH molecule. We investigate the energy landscape of RHF for both molecules with the varying bond lengths by searching for the constrained SPs of allowed indices. They are found by running Algorithm 2 from 1,000 random initial points (random seed $0\sim 999$); see the pseudocodes in Section 5.1 for initialization. For retraction, Algorithm 2 is equipped with the exponential mapping for the position part; the treatment for the direction part is similar to that in the previous subsection. The step size is specified as $\eta=10^{-1}$ and $\eta=10^{-2}$ for the H2 and LiH molecules, respectively, which are not necessarily optimal. The maximum iteration number and convergence tolerance are respectively set as $maxit=\infty$ and $tol=10^{-6}$. The results are compared with those obtained by solving FCI based on the RHF ground-state calculations, which is exact under the basis in use. In particular, FCI gives 16 states for the H2 molecule and 225 states for the LiH molecule.
@@
Figure 5: An overview of the constrained SPs of RHF found across the bond length interval on the H2 molecule. The blue solid and dashed lines represent the energies of FCI states belonging to the irreps $\rm A_{g}$} and $\rm A_{u}$}, respectively. The red dots, deepblue squares, yellow stars, and purple diamonds stand for the energies of the RHF SPs of indices 0, 1, 2, and 3, respectively. The zoom-in inset describes the FCI states by atomic fragments when the molecule is close to its dissociation limit.
@@
Figure 6: Zoom-in views of the constrained SPs of RHF. Upper left: an RHF index-1 SP disappears in (0.5 bohr, 0.6 bohr). Upper right: an RHF index-2 SP emerges in (1.6 bohr, 1.7 bohr). Lower: an RHF index-0 SP emerges in (2.6 bohr, 2.7 bohr).
Figure 7: An overview of the constrained SPs of RHF found across the bond length interval on the LiH molecule. The blue solid, dotted, and dashed lines represent the energies of FCI states belonging to the irreps A1, A2, B1 (or B2), respectively. The red dots, deepblue squares, yellow stars, purple diamonds, down-pointing cyan triangles, up-pointing orange triangles, left-pointing green triangles, right-pointing pink triangles, and navyblue pentagons stand for the energies of the RHF SPs of indices 0 to 8, respectively.
@@
Results on the H2 molecule. An overview of the constrained SPs of RHF identified across the considered bond length interval is shown in Figure 5, together with the ground-/excited-state energies of FCI as reference. It can be seen that the RHF energy landscape varies smoothly with the bond length. Note that FCI states are classified by their irreducible representations (irreps); for the H2 molecule in the computational point group $\rm D_{2h}$}, the relevant irreps are $\rm A_{g}$} and $\rm A_{u}$}, with eight FCI states belonging to each. It is observed that the RHF SPs are only able to describe a small subset (e.g., four near the equilibrium geometry, three close to the dissociation limit) of FCI states in terms of their energies. This behavior is consistent with the fact that RHF neglects electronic correlation [barca2014communication]. The deficiency can be mitigated to some extent by resorting to post-HF methods, which is investigated in a parallel work.
@@
Our results also reveal numerically that the varying bond length, as an external parameter, can lead to disappearance or emergence of constrained SPs. Zoom-in views are given in Figure 6. Concretely, in the interval (0.5 bohr, 0.6 bohr), an index-1 SP gets close to the index-2 SP, in terms of energy, and disappears; in the interval (1.6 bohr, 1.7 bohr), an index-2 SP close to the index-3 SP emerges; and in the interval (2.6 bohr, 2.7 bohr), an index-0 SP (i.e., a local minimizer) close to the index-1 SP emerges.
Results on the LiH molecule. The identified constrained SPs of RHF are shown in a similar way in Figure 7, again with the FCI states energies as reference. For the LiH molecule in the computational group $\rm C_{2v}$, the relevant irreps are A1, A2, B1, and B2. Note that the B1 and B2 FCI states are two-fold degenerate across the considered bond length interval. In this case, the RHF SPs appear to well describe the low-lying FCI states with energies lower than 6-6 hartree3331 hartree $\approx 4.36\times 10^{-18}$ J. But the FCI states with energies in (6-6 hartree, 4-4 hartree) are completely missed by the RHF SPs, probably due to inherent shortcomings of the RHF approximation. In addition, the potential energy surfaces of higher-index RHF SPs are found to be more complicated than those of low-index SPs, especially when the molecule is driven to the dissociation limit.
@@
The above results provide proof-of-concept evidence for the effectiveness of our algorithms in excited-state calculations. Nonetheless, we shall point out that a comprehensive quantum chemical analysis of the obtained SPs, though beyond the scope of present work, is essential for practical applications. As a nonlinear approximation to the exact theory, RHF may produce critical points that lack physical meanings; e.g., a spurious non-global local minimizer emerges when the bond length exceeds 2.7 bohr in the right panel of Figure 6. Moreover, since RHF is a low-dimensional approximation, a one-to-one correspondence between RHF SPs and excited states in the same energetic order no longer holds. Molecular geometries must also be taken into account. In the exact theory, the ESE is parameterized by atomic configurations. Variations in molecular geometry may induce degeneracies and crossings in states, which would lead to unfavorable features of potential energy surfaces within approximate theories (cf. Figure 6). From the computational aspect, it is of great importance to develop schemes capable of efficiently navigating the nonconvex landscape [xu2025general], instead of relying on random multi-start.
@@ Acknowledgements
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement EMC2 No 810367). The authors are grateful to Éric Cancès, Tony Lelièvre, and Panos Parpas for useful discussions.